OpenStudy (kidthatbro8):

What is 3^3/2 equal to? cube root of 9 square root of 9 cube root of 27 square root of 27

OpenStudy (anonymous):

\[\left( 3 \right)^{\frac{ 3 }{ 2 }}=\left[ \left( 3 \right)^3 \right]^{\frac{ 1 }{ 2 }}=\left( 27 \right)^{\frac{ 1 }{ 2 }} =?\]

OpenStudy (kidthatbro8):

i'm not really sure, @surjithayer

OpenStudy (anonymous):

\[x ^{\frac{ 1 }{ 2 }}=\sqrt{x}\]